OpenStudy (anonymous):

Pamela bought 3 t-shirts and 2 sweatshirts and her bill came to $73.50. Fulton bought 4 t-shirts and 1 sweatshirt for a total of $60.50. Olivia wants to buy 1 t-shirt and 1 sweatshirt. What will her total be?

OpenStudy (anonymous):

Please help.

OpenStudy (anonymous):

3x+2y=73.5 4x+1y=60.5 x=t_shirts y=sweatshirt

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

x=9.5 y=22.5

OpenStudy (anonymous):

x+y=32